evaluate without calculators or tables
Log 6 (one divided by 216) raised to power 2/3
Given that Log 2 to power 3 is 1.585 evaluated without using calculators
Log 2(24)
Log 8(24)

Answer by haileytucki(390)

log^6((1)/(216))^((2)/(3))
To find an approximate value of log^6((1)/(216))^((2)/(3)), use the formula log^na=(ln(a))/(ln(n)).
(ln((1)/(216)))/(ln(6))
Find the value of each expression using a calculator.
-3
Second Answer=
log^2(3)
To find an approximate value of log^2(3), use the formula log^na=(ln(a))/(ln(n)).
(ln(3))/(ln(2))
Find the value of each expression without using a calculator.
1.58



Third Answer=
log^2(24)
To find an approximate value of log^2(24), use the formula log^na=(ln(a))/(ln(n)).
(ln(24))/(ln(2))
Find the value of each expression using a calculator.
4.58


Fourth Answer=
log^8(24)
To find an approximate value of log^8(24), use the formula log^na=(ln(a))/(ln(n)).
(ln(24))/(ln(8))
Find the value of each expression using a calculator.
1.53